ganeshie8 (ganeshie8):

Are there infinitely many integers \(n\) such that both \(27n+23\) and \(169n+144\) are simultaneously prime ?

ganeshie8 (ganeshie8):

http://mathworld.wolfram.com/DirichletsTheorem.html

OpenStudy (ikram002p):

so just showing that gcd( 27n+23,169n+144 )=1 ?

OpenStudy (isaidavila):

Let's find out if 27 and 23 are relatively primes (27;23)=1 they are Now 144 and 169 and they are So now from dirichlet theorem there are infinite n numbers for this series contain infinite primes

OpenStudy (kainui):

My way that I think works is find a,b,x,y that make this possible for any integer k, \[27(ak+b)+23=169(xk+y)+144\] It ends up leading to basically solving this: \[27b=169y+121\] which has as one possible solution \[b=17\]\[y=2\] Anyways point is, now I can just pick \(n=ak+b=xk+y\) and by Dirichlet's theorem this should be just as good as any other solution...? I just gotta make sure stuff is relatively prime or something huh... I dunno I am too hyper from eating candy today to check my work sorry lol

OpenStudy (ikram002p):

@Kainui use Euclid algorithm to show they are relatively prime =D